Johnny is 10 and his granddad is 74. How many years ago was johnny's granddad 9 times as old as johnny?
Answer by ankor@dixie-net.com(22740)

Johnny is 10 and his granddad is 74.
How many years ago was johnny's granddad 9 times as old as johnny?
:
let y = no. of years ago that this was true
74 - y = 9(10 - y)
74 - y = 90 - 9y
-y + 9y = 90 - 74
8y = 16
y = 16/8
y = 2 yrs ago
:
:
Confirm this, 2 yrs ago grandpa was 72 and Johnny was 8. 8 * 9 = 72